{F} Find the solution of




(D^3+3D^2-4)y=0

1
Expert's answer
2022-02-20T08:06:14-0500

The auxiliary equation is "m^3+3m^2-4=0"

Performing synthetic division,


"\begin{array}{c|rrr}
1&1&3&0&-4\\
& & 1 & 4&4\\
\hline
&1&4&4&\mid 0

\end{array}"


We get "m-1" as a factor. Therefore,

"m^3+3m^2-4= (m-1)(m^2+4m+4)=0\\
\implies m=1~\&~ m^2+4m+4=0\\
i.e., m=1~\&~ m=-2(twice)\\"


The solution is "y = c_1 e^x + (c_2+c_3x)e^{-2x}"